The product of a number decreased by 3 and increased by 5 is 9. Find all numbers satisfying this condition.
A. Write the polynomial equation described above.
B. Solve the equation and enter your solutions below.
**I understand B I just have no idea how to do A, I missed class the other day and I am trying to get caught up before the next class.

Answer by stanbon(75887)

The product of a number decreased by 3 and increased by 5 is 9. Find all numbers satisfying this condition.
A. Write the polynomial equation described above.
------
Equation:: (x-3)(x+5) = 9
x^2 + 2x - 15 = 9

B. Solve the equation and enter your solutions below.
x^2 + 2x -24 = 0
-------
Factor:
(x+6)(x-4) = 0
x = -6 or x = 4
===================
Cheers,
Stan H.
===================